OpenStudy (anonymous):

write an equation of the line containing the specified point and perpendicular to the indicated line. (3, 1), 2x-3y=4

OpenStudy (anonymous):

1st find the slope of indicated line i.e 3y=2x-4 =>y=2/3x-4/3 here slope=2/3 but if the required line is perpndicular to indicate then ultiply slope by -1 i.e slope=m=-2/3 so now point slope form is (y-y1)=m(x-x1) (y-1)=-2/3(x-3) =>3(y-1)=-2(x-3) =?3y-3=-2x+6 =>3y+2x=9 answer

OpenStudy (anonymous):

get it..???

OpenStudy (anonymous):

Yes, thank you so much :)

OpenStudy (anonymous):

you welcome...

OpenStudy (anonymous):

You need to take the opposite reciprocal to get a perpendicular line, and not just the opposite @zaibali.qasmi?

OpenStudy (anonymous):

So m = -3/2 not -2/3